# Algebroid functions on complex manifolds

### v5

@@ Abstract.
We build a basic theory for algebroid functions on complex manifolds including several closely related topics such as uniformization problems, Nevanlinna theory, propagation of algebraic dependence and unicity problems, which is a development of the classical theory for algebroid functions on the complex plane. To the best of our knowledge, it is the first result for algebroid functions on a nontrivial complex manifold after H. Poincaré proposed the notion of algebroid functions.
@@
Algebroid functions; uniformization problems; second main theorem; Picard’s theorem; unicity.2010 Mathematics Subject Classification:
@@
- 3 Uniformization of algebroid functions on Hermitian manifolds
- 4 Nevanlinna theory of meromorphic functions on Kähler manifolds
- 5 Nevanlinna theory of algebroid functions on Kähler manifolds
- 6 Unicity problems of algebroid functions on Kähler manifolds
@@ 1.1. Background
where are holomorphic functions with no common zeros on An algebroid function is viewed as a special multi-valued analytic function, while which was first put forward, G. Darboux deemed that it is an important class of functions. To see that, for example, it is known that a -valued algebroid function defines a -leaf Riemann surface which is a (generic) -sheeted analytic covering of A further fact claims that can lift to a single-valued meromorphic function on via a natural projection satisfying that it therefore realizes the uniformization of In fact, it leads to the famous uniformization problem for Riemann surfaces, which had been solved by Poincaré-Klein-Koebe (see [28, 35]) who showed the uniformization theorem which asserts that every Riemann surface is conformally equivalent to some quotient space where is one of the three canonical domains: or (the unit disc), while is a proper discrete group of Möbius automorphisms of
Even so, the most concerned problem on algebroid functions is their value distribution. About 1930s after R. Nevanlinna [33] founded two fundamental theorems concerning the value distribution theory of meromorphic functions in 1925, many famous scholars such as G. Rémoundos, H. Selberg, E. Ullrich and G. Valiron, etc. began to study the value distribution theory of algebroid functions. For example, G. Rémoundos [36] proved a Picard’s theorem which asserts a nonconstant -valued algebroid function on can omit distinct values at most in Afterwards, G. Valiron [59] showed without details a five-value type theorem in unicity theory which claims that two nonconstant -valued algebroid functions on will be identical if they share +1 distinct values ignoring multiplicities in Since then, a large number of scholars have devoted themselves to this topic, involving investigations to Nevanlinna theory, Ahlfors theory of covering surfaces, unicity theory and normal family theory, etc.. More details refer to H. Cartan [9], He-Xiao [22], Niino-Ozawa [29], J. Noguchi [30], M. Ozawa [34], M. Ru [37, 38], Sun-Gao [41], J. Suzuki [43], L. Selberg [45, 46], N. Toda [54, 55], E. Ullrich [58] and their references therein.
The algebroid function theory in one complex variable has received a great deal of achievements, however, it didn’t mature until the last 30-40 years. So far, there is very little literature discussing the algebroid functions in several complex variables, except part of contributions of Hu-Yang [21] to the theory on due to some technical difficulties such as the complexity of branching points and multiple points as well as uniformization of an algebroid function. We even know nothing about the algebroid functions on a nontrivial complex manifold. Given two nonconstant algebroid functions on a non-compact Hermitian manifold (referred to (1) in Section 2.1 for definition). We are inspired by the following closely related fundamental problems:
@@
How many values do have to share at least to guarantee ?
To solve the first problem, we construct a multi-leaf complex manifold via which is a generic multi-sheeted analytic covering of Then, we lift to a meromorphic function on via the natural projection with The strategy of solving the second problem is to establish the Nevanlinna theory of algebroid functions on a complete Kähler manifold under certain curvature condition. In order to establish the Second Main Theorem of our trick used is to lift to via Thereout, one can receive the defect relation of which derives the Picard’s theorem. The central idea of dealing with the last problem is the applications of Nevanlinna theory. In doing so, we generalize for the first time the propagation theory of algebraic dependence introduced by L. Smiley [47] to the algebroid functions. This is a technical tool to study unicity problems. Based on it, we obtain the sharp lower bound of the number of the shared values of guaranteeing
@@ 1.2. Main results
Our study of algebroid function theory covers four closely related aspects: uniformization problems, Nevanlinna theory, propagation theory of algebraic dependence and unicity problems. So, we shall state the main results of this paper from the following four parts in turn.
A. Uniformization problems
Let be a non-compact Hermitian manifold of complex dimension At first, we generalize the notion of algebroid functions from to by means of some technical treatment, see Definition 2.15. Let be a -valued algebroid function on Assume that is a -leaf meromorphic function element of namely, is a local -valued meromorphic solution of equation (1) in a neighborhood of where is a branching point with order of If is a non-singular branching point of then we describe the asymptotic behavior of as approaches i.e., we show that can be expanded into the Pusieux series (see Theorem 2.14)
with around in a local holomorphic coordinate where are meromorphic functions in Pusieux series plays an important role in construction of the -leaf complex manifold of and so is further uniformized by lifting to That is:
Theorem A1.
@@ A. Uniformization problems
Theorem A1 can be applied directly to the extension of the fields of meromorphic functions. Let be a generic -sheeted analytic covering of We denote by the fields of meromorphic functions on respectively. As a consequence of Theorem A1, we deduce that
Corollary A2.
is a -degree algebraic extension of
@@ B. Nevanlinna theory
Nevanlinna theory in several complex variables has been developed rapidly, due to the outstanding work of a number of excellent scholars. For example, referring to Bott-Chern [8], Carlson-Griffiths-King [10, 20], H. Fujimoto [17], K. Kodaira [25], Noguchi-Winkelmann [31, 32], M. Ru [38], B. Shiffman [40], F. Sakai [42], W. Stoll [48, 49, 52] and Y. T. Siu [53], etc., and also refer the reader to A. Atsuji [1, 2, 3] and X. J. Dong [14, 15], etc.. In this paper, our central work is to establish the Nevanlinna theory of algebroid functions on a non-compact complete Kähler manifold based on estimates of Li-Yau [27] on heat kernels and so we can use which to derive the Picard’s theorem.
In this part and the following Parts B, C, D, is a non-compact complete Kähler manifold. Let be a -valued algebroid function on See definition in Section 5.3, the Nevanlinna’s functions of are defined under two kinds of different curvature conditions. To unify the form, we write
(b1) has non-negative Ricci curvature and maximal volume growth
Refer to [44, 56, 57], we see that such manifolds exist under this curvature condition. Let be the Ricci form of then we obtain:
@@
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
@@
We obtain the following Picard’s theorem:
Corollary B3 (Picard Theorem).
Let be a transcendental -valued algebroid function on Then, can omit distinct values at most in
@@
Let be the curvature function defined by (24) in Section 4.3. Under this curvature condition, we obtain:
@@
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
@@
Assume the same conditions as in Theorem If satisfies the growth condition
@@
We obtain the following Picard’s theorem:
Corollary B6 (Picard Theorem).
Let be a nonconstant -valued algebroid function on satisfying the growth condition
Then, can omit distinct values at most in
C. Propagation theory of algebraic dependence
Propagation theory of algebraic dependence was introduced by L. Smiley [47]. Since it has important applications in unicity problems, then it aroused widespread concern. For example, the reader may refer to Y. Aihara [5, 6, 7], Dulock-Ru [12], S. Drouilhet [13], H. Fujimoto [18, 19], S. Ji [23, 24] and W. Stoll [51], etc.. The paper extends this theory to the algebroid functions via the Nevanlinna theory established in Part B.
@@ C. Propagation theory of algebraic dependence
In addition, if has non-positive sectional curvature, then we ask for
@@ D. Unicity problems
Five-value theorem is a well-known theorem in unicity theory, which states that two nonconstant meromorphic functions on are identical if they share five distinct values (ignoring multiplicities) in This theorem was later extended to algebroid functions on by G. Valiron [59] and further improved by Sun-Gao [41]. In our study, a sharp five-value type theorem for algebroid functions on a non-compact complete Kähler manifold is proved by applying the propagation theory of algebraic dependence established in Part C.
(d1) has non-negative Ricci curvature and maximal volume growth
@@
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that for all If satisfies
with then
@@
Let be nonconstant -valued, -valued algebroid functions on respectively satisfying the growth condition
Let be distinct values in Assume that for all If satisfies
with then
@@
Let be nonconstant -valued, -valued algebroid functions on respectively, satisfying the growth condition
If share distinct values ignoring multiplicities in then
@@ 2.1. Notion of algebroid functions
First of all, we shall illustrate how the notion of algebroid functions on can be generalized to non-compact Hermitian manifolds. This is the starting point of a theory on algebroid functions we would like to build in this paper.
Let be a -dimensional non-compact Hermitian manifold. We consider an irreducible algebraic equation of the unknown function over :
where are holomorphic functions locally defined on such that they define a meromorphic mapping
@@
The local definition for means that a collection of holomorphic functions i.e., is a holomorphic function for every satisfying that
@@ 2.1.1. Meromorphic function elements
Evidently, we see that is continuous on
Next, we show that is holomorphic on One just needs to show without loss of generality that is analytic at Taking a local holomorphic coordinate around such that Apply Hartog’s theorem, it is sufficient to show that is holomorphic in every holomorphic direction at For and , set
@@
To see the common roots of and we define the resultant of and by a determinant:
@@
If then is a multiple root of as a polynomial in If we set and rewrite (1) as where
@@ 2.1.2. Analytic continuation of meromorphic function elements
We equip with a topology. A meromorphic function element is said to be a direct analytic continuation of a meromorphic function element if and in a neighborhood of Apply analytic continuation, we can regard that
@@
Pick If then we can consider -neighborhoods of and of with Evidently, we have If for then we can take a function element It implies that in a small neighborhood of By the uniqueness theorem of analytic functions, we obtain but it contradicts with ∎
Next, we introduce the notion of analytic continuation (by which we mean both holomorphic continuation and meromorphic continuation of meromorphic function elements.
@@
The analytic continuation of elements in along a curve in is a continuous mapping of into namely, an analytic continuation of to along is a curve connecting and Indeed, if is an analytic continuation of along a curve, then is also an analytic continuation of along the same curve. Thus, the analytic continuation is symmetric.
@@
Let be distinct function elements at every fixed With the aid of Corollary 2.11, then we only need to prove that all can continue analytically to each other. Let be the largest integer such that there are function elements in that can continue analytically to each other along a curve. Clearly, we have Without loss of generality, one may assume that these function elements are So, we can write
@@
It is equivalent to show that: can continue analytically to the whole Apply Theorem 2.10, can continue analytically to along a curve at any where is a permutation of By Vieta’s theorem, determine the same in a small neighborhood of i.e., continue analytically to Hence, holds. If then we are able to determine another set of using the rest function elements However, this implies that equation (1) is reducible, a contradiction. ∎
@@ 2.1.3. Properties of meromorphic function elements near multiple points
We discuss the properties of function elements near multiple points. Let be a pair satisfying that and
Take a sufficiently small neighborhood of such that on By the continuation of there exists a number such that on Let then there exist distinct regular function elements and in respectively. Make use of Theorem 2.10, then for a curve with and are able to continue analytically to along In other words, there exists a permutation of such that can continue analytically to along respectively. By symmetrical property, can also continue analytically to along However, will not, in general, continue analytically back to along respectively. In the following, we will investigate the analytic continuation of a function element along a closed Jordan curve. Let and pick as a closed Jordan curve passing by around Without loss of generality, we extend analytically along started at When returns to we attain to a function If then one continues to extend analytically along Repeat this action, then we will see that cycles on and on periodically. Let be the smallest period such that returns to itself, we obtain We say that is a branching point with order (or multiplicity) of if Repeatedly, then is divided into finite groups. Let the number of groups. If the -th group has members, then we have
Without loss of generality, we consider the first group: Take a proper there is a biholomorphic mapping such that
@@
Set and write Next, we shall show that is a single-valued analytic function on where is the unit disc centered at in the -plane. Evidently, is holomorphic on We still need to prove that is a removed singularity of In doing so, it is sufficient to show that is bounded on due to Riemann’s analytic continuation. There exists a number such that on for Combining (1) with on we obtain
@@
According to the above arguments, we show that is a -valued analytic function near In other words, if we cut along a (real) hypersurface which passes through then is separated into single-valued analytic components, and these components replace each other when they continue analytically along a curve around in Observe the local expansion of is a -multiple -valued branching point (with order ) of
Assume that is not a singular branching point of Set we write
@@
Thus, i.e., the case turns to the case Note that is a pole of order of
To conclude, for each we have confirmed that there are distinct single-valued meromorphic function elements subordinate to equation (1) if is not a branching point of If is a branching point of then we can have several multi-valued function elements subordinate to equation (1), say with where gives a branching point of with order In particular, if is a non-singular branching point of then there is the Pusieux expansion
near in a local holomorphic coordinate Note that is a -multiple root of namely, can separate into single-valued components with value at If then can separate into single-valued components which take at In particular, is a meromorphic function element for Furthermore, is a regular function element if also. By analytic continuation, equation (1) uniquely determines a -valued meromorphic function on outside a set of branching points of What is more, is separated into single-valued meromorphic components on a cut-manifold, say where is any real hypersurface of passing by all the branching points of such that is connected. Let denote the set of all the branching points of which is an analytic set contained in called the branching set of We obtain:
@@
in a local holomorphic coordinate near such that and for where is the branching set in a local neighborhood of and are meromorphic functions in
Now, we introduce the definition of algebroid functions on a non-compact Hermitian manifold as follows.
@@
where are holomorphic functions locally defined on such that they define a meromorphic mapping from into
Remark 2.16.
By Definition 2.15, we see that the condition for is also necessary due to Vieta’s theorem. In other words, a -valued algebroid function on is defined equivalently by an irreducible algebraic equation
where are meromorphic functions on

### v9

# Nevanlinna Theory of Algebroid Functions on Complete Kähler Manifolds

@@ Abstract.
We generalize the classical Nevanlinna theory for algebroid functions on to a complete Kähler manifold with non-negative Ricci or non-positive sectional curvature. As an application, we give some Picard type theorems and five-value type theorems for algebroid functions under certain conditions.
@@
algebroid functions; Nevanlinna theory; second main theorem; Picard type theorem.2010 Mathematics Subject Classification:
@@
- 3 Uniformization of Algebroid Functions
- 4 Nevanlinna Theory of Meromorphic Functions
- 5 Nevanlinna Theory of Algebroid Functions
- 6 Unicity Problems of Algebroid Functions
@@ 1.1. Motivation
where are holomorphic functions with no common zeros on An algebroid function is viewed as a special multi-valued analytic function, while which was first put forward, G. Darboux deemed that it is an important class of functions. To see that, for example, it is known that a -valued algebroid function defines a -leaf Riemann surface which is a (generic) -sheeted analytic covering of A further fact claims that can lift to a single-valued meromorphic function on via a natural projection such that which is called the uniformization of
The most concerned problem on algebroid functions seems to be the value distribution. In the 1930s after R. Nevanlinna [22] founded two fundamental theorems concerning the value distribution theory of meromorphic functions in 1925, many famous scholars such as G. Rémoundos, H. Selberg, E. Ullrich and G. Valiron, etc. began to study the value distribution theory of algebroid functions. For example, G. Rémoundos [24] proved a Picard’s theorem which claims that a nonconstant -valued algebroid function on omits distinct values at most in Afterwards, G. Valiron [40] showed without details a five-value type theorem in unicity theory, which claims that two nonconstant -valued algebroid functions on will be identical if they share +1 distinct values ignoring multiplicities in Since then, a large number of scholars have devoted themselves to this topic, including investigations to Nevanlinna theory, Ahlfors theory of covering surfaces, unicity theory and normal family theory, etc. More details refer to H. Cartan [7], He-Xiao [16], Niino-Ozawa [19], J. Noguchi [20], M. Ozawa [23], M. Ru [25, 26], Sun-Gao [28], J. Suzuki [29], L. Selberg [30, 31], N. Toda [37, 38] and E. Ullrich [39], etc.
Algebroid function theory in one complex variable didn’t mature until the last 30-40 years. However, there is little literature on the algebroid functions in several complex variables, except for the work of Hu-Yang [15]. Given two nonconstant algebroid functions on a non-compact Hermitian (Kähler) manifold . We are inspired by the following three fundamental problems:
@@
How many values should share at least to guarantee ?
In this paper, we answer these three problems by generalizing the classical Nevanlinna theory for algebroid functions on to a non-compact complete Kähler manifold with non-negative Ricci or non-positive sectional curvature.
@@ 1.2. Main Results
A. Uniformization
Let be a non-compact Hermitian manifold of complex dimension At first, we generalize the notion of algebroid functions from to by means of a technical treatment (see Definition 2.15). Let be a -valued algebroid function on Assume that is a -leaf meromorphic function element of namely, is a local -valued meromorphic solution of equation (7) in a neighborhood of where is a branching point with order of If is a “non-singular” branching point of then we describe the asymptotic behavior of as approaches i.e., we prove that can be expanded into the Pusieux series (see Theorem 2.14)
with around in local holomorphic coordinates where are meromorphic functions in With this Pusieux series, we show that
Theorem A.
@@ B. Nevanlinna theory
In this part and the following Parts C, D, we utilize to stand for a non-compact complete Kähler manifold. Let be a -valued algebroid function on Refer to the definition in Section 5.3, we have the Nevanlinna’s functions of and on where is the Ricci form of and is a precompact domain containing a fixed reference point To unify the form, we write
The simple defect of with respect to is defined by
(b1) has non-negative Ricci curvature
Assume that is non-parabolic. Set
where denotes the Riemannian volume of the geodesic ball centered at with radius
@@
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is defined by
When we have Hence, Theorem B1 covers the classical second main theorem for with (see Corollary 5.5).
@@
if one of the following conditions is satisfied
satisfies the volume growth condition
where is defined by
is of non-polynomial type growth, i.e., satisfies the growth condition
We have as due to
The condition (2) is relaxed. For example, it holds when with and holds when with or and
Corollary B3 (Picard Type Theorem).
Let be a nonconstant -valued algebroid function on Then, can omit distinct values at most in if one of the conditions in Corollary is satisfied.
@@
Assume that is non-positively curved satisfying
for constants with that either or where are the sectional curvature and Ricci curvature of respectively. Set
@@
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is defined by
Theorem B4 covers the classical second main theorem for both and the unit ball with Poincaré metric in (see Corollaries 5.7 and 5.8).
@@
Assume the same conditions as in Theorem B4. If
@@
Corollary B6 (Picard Type Theorem).
Let be a nonconstant -valued algebroid function on satisfying Then, can omit distinct values at most in
C. Propagation of algebraic dependence
Propagation problem for algebraic dependence of meromorphic mappings was first studied by L. Smiley [32] and further studied by Y. Aihara [3, 4, 5], S. Drouilhet [9], H. Fujimoto [13, 14], Dulock-Ru [10] and W. Stoll [36], etc.
Assume that is either non-parabolic with non-negative Ricci curvature, or has non-positive sectional curvature satisfying (3).
@@ C. Propagation of algebraic dependence
and (2) is satisfied if is non-parabolic with non-negative Ricci curvature, or (5) is satisfied if has non-positive sectional curvature satisfying (3).
@@ D. Unicity problems
(d1) has non-negative Ricci curvature
Assume that is non-parabolic satisfying (2). Let be a positive integer. Consider the following condition
with
@@
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that for all If satisfies then
@@
Let be nonconstant -valued, -valued algebroid functions on satisfying respectively. Let be distinct values in Assume that for all If satisfies then
@@
Let be nonconstant -valued, -valued algebroid functions on satisfying respectively. If share distinct values ignoring multiplicities in then
@@ 2.1. Notion of Algebroid Functions
First of all, we would like to extend the notion of algebroid functions to a non-compact Hermitian manifold. Let be a -dimensional non-compact Hermitian manifold. Consider an irreducible algebraic equation in on :
where are holomorphic functions locally defined on such that they well define a meromorphic mapping
@@
The local definition for means that a collection of holomorphic functions i.e., is a holomorphic function for all satisfying that
@@ 2.1.1. Meromorphic function elements
Evidently, is continuous on
Next, we show that is holomorphic on One just needs to show without loss of generality that is analytic at Taking local holomorphic coordinates around such that Apply Hartog’s theorem, it is sufficient to show that is holomorphic in every holomorphic direction at For and , set
@@
To see the common roots of and we define the resultant of and by a -order determinant:
@@
If then has a multiple root as a polynomial in If we set and rewrite (7) as where
@@ 2.1.2. Analytic continuation of meromorphic function elements
We equip with a topology. A meromorphic function element is said to be a direct analytic continuation of a meromorphic function element if and in a neighborhood of By the analytic continuation, we can regard that
@@
Pick If then we can consider -neighborhoods of and of where Evidently, we have If for then we can take a function element It implies that in a small neighborhood of By the uniqueness theorem of analytic functions, we obtain but it contradicts with ∎
Next, we introduce the notion of analytic continuation (by which we mean both holomorphic continuation and meromorphic continuation) of meromorphic function elements.
@@
The analytic continuation of elements in along a curve in is a continuous mapping of into namely, an analytic continuation of to along is a curve connecting and Moreover, if is an analytic continuation of along a curve, then is also an analytic continuation of along the same curve. Thus, the analytic continuation is symmetric.
@@
Let be distinct function elements at every fixed With the aid of Corollary 2.11, then we only need to prove that all can continue analytically to each other. Let be the largest integer such that there are function elements in which can continue analytically to each other along a curve. Clearly, we have Without loss of generality, we shall assume that these function elements are So, we can write
@@
It is equivalent to show that: can continue analytically to the whole Apply Theorem 2.10, can continue analytically to along a curve at any where is a permutation of By Vieta’s theorem, determine the same in a small neighborhood of i.e., continue analytically to Hence, holds. If then we are able to determine another set via the rest function elements However, it implies that equation (7) is reducible, which is a contradiction. ∎
@@ 2.1.3. Properties of meromorphic function elements near multiple points
We discuss the local properties of function elements near multiple points. Let be a pair satisfying that and
Taking a sufficiently small neighborhood of such that on By the continuation of there exists a number such that on Let then there exist distinct regular function elements and in respectively. By the use of Theorem 2.10, then for a curve with and are able to continue analytically to along In other words, there exists a permutation of such that can continue analytically to along respectively. By symmetrical property, can also continue analytically to along However, will not, in general, continue analytically back to along respectively. In the following, we discuss the analytic continuation of a function element along a closed Jordan curve surrounding a multiple point. Take Pick as a closed Jordan curve passing by around Without loss of generality, we only extend analytically along starting from When returns to from along one will attain to a function If then one continues to extend analytically along Repeating this action, we see finally that cycles on and on periodically. Let be the smallest period such that returns to itself, we obtain We say that is a branching point with order (multiplicity) of if Repeatedly, then it is not hard to verify that can be divided into several groups. Let be the number of groups. If the -th group has members, then we have
Without loss of generality, we consider the first group: Take a proper we have a biholomorphic mapping such that
@@
Set and write Next, we shall show that is a single-valued analytic function on where is the unit disc centered at in the -plane. Evidently, is holomorphic on We still need to prove that is a removed singularity of In doing so, it is sufficient to show that is bounded on due to Riemann’s analytic continuation. There exists a number such that on for Combining (7) with that on we obtain
@@
According to the above arguments, we show that is a -valued analytic function near In other words, if we cut along a (real) hypersurface which passes through then is separated into single-valued analytic components, and these components replace each other when they continue analytically along a closed curve around in By a local expansion of is a -multiple -valued branching point (with order ) of
Assume that is not a singular branching point of Set and write
@@
Thus, i.e., the case turns to the case Note that is a pole with order of
To conclude, for each we have confirmed that there are distinct (single-valued) meromorphic function elements subordinate to equation (7), if is not a branching point. If is a branching point, then one shows that there are several multi-valued function elements (which contain single-valued function elements) subordinate to equation (7), say where is a -order branching point of satisfying At this time, we call a -order branching point of If is a non-singular branching point of then we have the convergent Pusieux expansion
near in a local holomorphic coordinate Note that is a -multiple root of namely, can separate into single-valued components with value at If then can separate into single-valued components which take at In particular, is a meromorphic function element for Furthermore, is a regular function element if also. By the analytic continuation, equation (7) uniquely defines a -valued meromorphic function on outside a set of branching points of and is separated into single-valued meromorphic components on Let denote the set of all branching points of which is an analytic set contained in called the branching set of As a result, we conclude that
@@
in a local holomorphic coordinate near such that =0 and for where is the branching set in a local neighborhood of and are meromorphic functions in
We conclude the definition of algebroid functions on as follows.
@@
where are holomorphic functions locally defined on such that they well define a meromorphic mapping from into
Let denote distinct (single-valued) meromorphic components of Given any -leaf function element of with If denote the single-valued components of then there exists a permutation of such that on In other words, defines a -leaf function element on Thus, any function element of can be represented locally by one or several single-valued components of